# InFL-UX: A Toolkit for Web-Based Interactive Federated Learning

### v1

@@ Abstract.
This paper presents InFL-UX, an interactive, proof-of-concept browser-based Federated Learning (FL) toolkit designed to integrate user contributions seamlessly into the machine learning (ML) workflow. InFL-UX enables users across multiple devices to upload datasets, define classes, and collaboratively train classification models directly in the browser using modern web technologies. Unlike traditional FL toolkits, which often focus on backend simulations, InFL-UX provides a simple user interface for researchers to explore how users interact with and contribute to FL systems in real-world, interactive settings. By prioritising usability and decentralised model training, InFL-UX bridges the gap between FL and Interactive Machine Learning (IML), empowering non-technical users to actively participate in ML classification tasks.
@@ 1. Introduction and Related Work
Traditional machine learning (ML) often struggles with limited data, particularly in specialised domains where data collection is costly or labour-intensive (Roh et al., 2021). The lack of direct involvement of domain experts, due to technical barriers, further delays the acquisition of new training data (Holstein, 2024). To address these challenges, Fails and Olsen (2003) introduced interactive machine learning (IML), enabling non-technical users to train ML models using their own data through manual classification or correcting model outputs. Unlike traditional ML, IML allows real-time updates in response to user input, facilitating focused and incremental adjustments (Amershi et al., 2014; Dudley and Kristensson, 2018). Building on these advancements, Tseng et al. (2023) developed Co-ML, a tablet-based application for collaboratively building ML image classification models across multiple devices, focusing on teaching dataset design practices by creating a shared dataset. In this paper, we extend these concepts by proposing a browser-based tool that allows users to collaborate on IML tasks using federated learning.
@@
Typically, ML models are trained on centralised datasets. However, in domains such as healthcare, data is often distributed across multiple devices and cannot be shared due to privacy concerns. Federated learning (FL) addresses this limitation by enabling decentralised training of a shared model while keeping data on the clients’ devices (McMahan et al., 2017). FL operates by aggregating locally computed updates to a central server. Two key types of FL algorithms exist: synchronous and asynchronous. Synchronous FL (McMahan et al., 2017) requires all participating clients to train simultaneously after receiving the latest global model, submitting updates together for incorporation into the global model. In contrast, asynchronous FL (Xie et al., 2019) allows clients to train and submit updates independently, with the server updating the global model upon each arrival. Given its flexibility, asynchronous FL is more suitable for our use case. Recent advancements in JavaScript-based deep learning (DL) frameworks, such as TensorFlow.js (Smilkov et al., 2019), have made browser-based DL feasible (Ma et al., 2019). Google’s Teachable Machine (Carney et al., 2020) demonstrates this with a no-code web application that enables users to train ML models locally via an intuitive interface. However, this tool focuses exclusively on local training without integrating FL capabilities. In contrast, browser-based FL frameworks, such as those by Lian et al. (2022) and Ángel Morell and Alba (2022), support FL but lack interactivity, requiring users to provide data in predefined folders with minimal engagement.
@@
In this paper, we introduce InFL-UX111The name is a combination of Interactive Machine Learning, Federated Learning, and User Experience., an interactive, browser-based FL toolkit that demonstrates a proof-of-concept (POC) implementation leveraging modern web technologies. The application employs asynchronous FL, implemented using FedAsync (Xie et al., 2019), a method chosen for its simplicity and single-hyperparameter configuration, enabling collaborative model training. Users can upload datasets, define custom classes, and train ML models tailored to diverse classification tasks. While most FL research focuses on simulations, InFL-UX combines FL and IML, prioritising user engagement to address FL challenges from a user-centric perspective. Our goal is to assist FL practitioners in integrating FL setups into intelligent user interfaces and evaluating them in user studies. Future development will enhance the system with advanced FL aggregation strategies and ML tasks, such as data annotation, to increase adoption across various domains.
@@ 2. System Design
InFL-UX utilises ONNX Runtime222https://onnxruntime.ai/ (Accessed January 03, 2025), a cross-platform machine learning library designed for fast on-device inference and training, with support for web browsers. It leverages modern browser APIs such as WebAssembly (Rossberg, 2022), a low-level, assembly-like language enabling near-native performance for web applications, and WebGPU (Ninomiya et al., 2024), which allows high-performance GPU computations and improves compatibility with modern GPUs as the successor to WebGL333https://www.khronos.org/webgl/ (Accessed January 03, 2025). This browser-based implementation facilitates easy integration with existing systems due to the ubiquity of web browsers and offers seamless compatibility with models from popular ML frameworks through the ONNX (Open Neural Network Exchange) format. The application is developed in Python using Flask444https://flask.palletsprojects.com/en/stable/ (Accessed January 03, 2025), and integrates ONNX Runtime via JavaScript and encapsulates UI components using Web Components555https://www.webcomponents.org/ (Accessed January 03, 2025), ensuring independence from specific UI frameworks. Persistent browser storage is achieved through IndexedDB (Bell, 2024), enabling the retention of user data, such as uploaded files and inference results, between sessions, even after the browser is closed. The code for InFL-UX is publicly available online: https://github.com/tmaurer42/interactive-fl-poc.
@@ 2.1. System and Application Design
The system architecture, depicted in Figure 1, consists of two main components: the FL Server and the FL Web Server. The FL Web Server hosts the web application and delivers necessary files to client devices, while the FL Server manages client updates, maintains the global model, and provides the latest model version and training instructions upon request. This architecture underpins the POC application, which focuses on image classification but can also support tasks like object detection and image segmentation via abstract classes. The system enables the simulation of multiple independent clients to demonstrate its functionality. One client acts as the administrator, configuring the setup through the configuration page (Figure 2a) by defining the use case, selecting the aggregator and ML model, and setting training parameters. On the client side, users upload images, receive suggested labels based on the current model, and can accept or modify these labels (Figure 2b). After reviewing, a training session can be initiated using the local dataset (Figure 2c). Upon training completion, updates are sent automatically to the central server for aggregation. Clients can also test the global model using their own testing datasets (Figure 2d).
Currently, the system supports a single FL aggregation strategy, FedAsync, and one ML model, MobileNetV2 (Sandler et al., 2019). Extending the system to include a new aggregation strategy or ML model is straightforward and requires minimal code changes. A new aggregation strategy involves defining the strategy as a Python function and adding its configurable parameters to the administrator’s configuration interface. Similarly, new ML models can be incorporated by extending the administrator interface’s dropdown menu with additional PyTorch models666https://pytorch.org/ (Accessed January 03, 2025). The system is deployed using Docker777https://www.docker.com/ (Accessed January 03, 2025). InFL-UX is designed to be extensible, enabling the integration of additional aggregation strategies. However, it does not support simulation tasks, as these are handled by established frameworks such as Flower (Beutel et al., 2022). Instead, InFL-UX serves as a subsequent step following the simulation phase, focusing on validating aggregation methods in a real-world context.
@@ 2.2. Limitations
Several limitations emerged during the development of the application. The ONNX Runtime integration for the web, being relatively new, has incomplete functionalities for model training. For instance, the learning rate of the optimiser cannot be adjusted, requiring clients to use a default rate of 0.001, and only a limited set of loss functions are available. Additionally, WebGPU currently supports only inference, while training is restricted to the WebAssembly backend. Browser-specific constraints also affect the application. The size of the local dataset stored in the browser’s IndexedDB is restricted, with storage limits varying by browser and disk space availability. Furthermore, the WebGPU API is currently accessible only in developer builds of modern browsers, which are often disallowed on systems with strict security policies.
@@ 3. Conclusion and Future Work
The development of the POC InFL-UX demonstrated the feasibility of implementing an interactive, browser-based FL system, integrating state-of-the-art technologies from the UI to the server-side aggregation process. The approach opens the door for integrating FL capabilities into IML systems, especially in privacy-sensitive domains. Despite its limitations, the application successfully utilised ONNX Runtime and modern browser features, such as WebAssembly and WebGPU, to enable client-side training for FL. The framework supports collaborative training across devices while maintaining compatibility and promoting user adoption.
Future work should focus on conducting user studies to evaluate adoption and performance in real-world scenarios. Expanding the application to include additional computer vision tasks, such as image segmentation or object detection, is another promising direction. Integrating explainable AI techniques, such as class activation maps (Selvaraju et al., 2017), could enhance transparency and user trust. Additionally, enabling administrators to deploy new training tasks and allowing clients to participate in multiple tasks with distinct datasets for each use case would significantly broaden the application’s utility.
@@ Acknowledgements.
This work was funded, in part, by the German Federal Ministry of Education and Research (BMBF) under grant number 01IW23002 (No-IDLE).References
@@
- Carney et al. (2020) Michelle Carney, Barron Webster, Irene Alvarado, Kyle Phillips, Noura Howell, Jordan Griffith, Jonas Jongejan, Amit Pitaru, and Alexander Chen. 2020. Teachable Machine: Approachable Web-Based Tool for Exploring Machine Learning Classification. In Extended Abstracts of the 2020 CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿Honolulu¡/city¿, ¡state¿HI¡/state¿, ¡country¿USA¡/country¿, ¡/conf-loc¿) (CHI EA ’20). Association for Computing Machinery, New York, NY, USA, 1–8. https://doi.org/10.1145/3334480.3382839

### v2

@@ Abstract.
This paper presents InFL-UX, an interactive, proof-of-concept browser-based Federated Learning (FL) toolkit designed to integrate user contributions into the machine learning (ML) workflow. InFL-UX enables users across multiple devices to upload datasets, define classes, and collaboratively train classification models directly in the browser using modern web technologies. Unlike traditional FL toolkits, which often focus on backend simulations, InFL-UX provides a simple user interface for researchers to explore how users interact with and contribute to FL systems in real-world, interactive settings. InFL-UX bridges the gap between FL and interactive ML by prioritising usability and decentralised model training, empowering non-technical users to actively participate in ML classification tasks.
@@ 1. Introduction and Related Work
Traditional machine learning (ML) is often constrained by limited data, particularly in specialised domains where data acquisition is expensive or labour-intensive (Roh et al., 2021). Moreover, technical barriers hinder direct input from domain experts, further delaying new data collection (Holstein, 2024). To overcome these issues, Fails and Olsen (2003) introduced interactive machine learning (IML), which enables non-technical users to train ML models by manually classifying data or correcting outputs. Unlike conventional ML, IML supports real-time updates based on user input, permitting focused, incremental refinements (Amershi et al., 2014; Dudley and Kristensson, 2018; Tusfiqur et al., 2022). Extending this work, Tseng et al. (2023) developed Co-ML, a tablet-based application for collaboratively constructing ML image classification models that emphasises shared dataset design. In this paper, we build on these concepts by proposing a browser-based tool that facilitates collaborative IML using federated learning (FL).
@@
ML models are traditionally trained on centralised datasets. However, in fields such as healthcare, data are distributed across multiple devices and cannot be shared due to privacy constraints. FL mitigates this issue by enabling decentralised training of a shared model while retaining data on client devices (McMahan et al., 2017). FL aggregates local updates on a central server and comes in two main forms: synchronous and asynchronous. Synchronous FL (McMahan et al., 2017) requires all clients to train concurrently with the latest global model and submit updates together. In contrast, asynchronous FL (Xie et al., 2019) allows clients to train and submit updates independently, with the server updating the global model upon each submission. Due to its flexibility, asynchronous FL is more suited to our application. Recent advances in JavaScript-based deep learning frameworks, such as TensorFlow.js (Smilkov et al., 2019; Ma et al., 2019), have made browser-based deep learning (DL) feasible. For example, Google’s Teachable Machine offers a no-code interface for local model training; however, it is limited to local training and does not incorporate FL. In contrast, browser-based FL frameworks proposed by Lian et al. (2022) and Ángel Morell and Alba (2022) support FL but lack interactive elements, requiring users to supply data in predefined folders.
@@
In this work, we introduce InFL-UX111The name is a combination of Interactive Machine Learning, Federated Learning, and User Experience., an interactive, browser-based FL toolkit that demonstrates a proof-of-concept (POC) implementation using contemporary web technologies. The application employs asynchronous FL via FedAsync (Xie et al., 2019), chosen for its simplicity and single-hyperparameter configuration, to enable collaborative model training. Users can upload datasets, define custom classes, and train models for various classification tasks. While much FL research relies on simulations, InFL-UX integrates FL with IML to prioritise user engagement and address FL challenges from a user-centric perspective. We aim to assist FL practitioners in embedding FL into intelligent user interfaces (UIs) and evaluate these setups through user studies. Future enhancements will incorporate advanced FL aggregation methods and additional ML tasks, such as data annotation, to broaden applicability.
@@ 2. System Design
InFL-UX utilises ONNX Runtime222https://onnxruntime.ai/ (Accessed January 03, 2025), a cross-platform ML library that supports fast on-device inference and training in web browsers. It leverages modern browser APIs, including WebAssembly (Rossberg, 2022)—a low-level language offering near-native performance—and WebGPU (Ninomiya et al., 2024), which facilitates high-performance GPU computations and supersedes WebGL333https://www.khronos.org/webgl/ (Accessed January 03, 2025). This browser-based approach, which capitalises on the ubiquity of web browsers, ensures seamless compatibility with models from prevalent ML frameworks via the ONNX format. Developed in Python using Flask444https://flask.palletsprojects.com/en/stable/ (Accessed January 03, 2025), the application integrates ONNX Runtime through JavaScript and employs Web Components555https://www.webcomponents.org/ (Accessed January 03, 2025) for UI modularity, thereby avoiding dependence on specific UI frameworks. Persistent storage is managed through IndexedDB (Bell, 2024), allowing user data, such as uploaded files and inference results, to be retained across sessions. The code for InFL-UX is publicly available at https://github.com/tmaurer42/interactive-fl-poc.
This figure shows the overall architecture of the application, structured in a client-server setup.
@@ 2.1. System and Application Design
The architecture (Figure 1) comprises two main components: the FL Server and the FL Web Server. The FL Web Server hosts the web application and delivers required files to client devices, while the FL Server manages client updates, maintains the global model, and provides the latest model version and training instructions on demand. This structure underpins the POC, which primarily addresses image classification; it can also accommodate tasks such as object detection and image segmentation through abstract classes and supports the simulation of multiple independent clients. An administrator client configures the system via the configuration page (Figure 2a) by specifying the use case, selecting the aggregator and ML model, and setting training parameters. Clients then upload images, receive model-generated label suggestions, and may accept or adjust these labels (Figure 2b). Once reviewed, a training session is initiated locally, and upon completion, updates are automatically forwarded to the central server for aggregation. Clients can evaluate the global model using separate testing datasets (Figure 2c).
This figure displays multiple screenshots from the InFL-UX application, highlighting various user interactions. Starting with the training configuration window in (a), then the label review process for the training images, which allows users to correct or accept the suggested labels in (b), and lastly, the interface for testing the trained model shown in (c).
Currently, the system implements a single FL aggregation strategy, i.e., FedAsync, and one ML model, i.e., MobileNetV2 (Sandler et al., 2019). Adding new aggregation strategies or ML models is straightforward, requiring only minimal code modifications: a new aggregation strategy is defined as a Python function with corresponding configurable parameters added to the administrator interface, and additional ML models are integrated via an expanded dropdown menu featuring extra PyTorch models666https://pytorch.org/ (Accessed January 03, 2025). The system is deployed using Docker777https://www.docker.com/ (Accessed January 03, 2025). Although designed for extensibility, InFL-UX does not support simulation tasks, which are handled by established frameworks such as Flower (Beutel et al., 2022); instead, it focuses on validating aggregation methods in real-world settings.
@@ 2.2. Limitations
Several limitations emerged during development. The relatively new ONNX Runtime integration for the web offers incomplete training functionalities; for example, the optimiser’s learning rate is fixed at 0.001, and only a limited set of loss functions is available. Furthermore, while WebGPU currently supports inference, training is restricted to the WebAssembly backend. Browser-specific constraints also affect the application: the local dataset stored in IndexedDB is limited in size (varying by browser and available disk space), and the WebGPU API is presently available only in developer builds of modern browsers, which may be restricted by security policies.
@@ 3. Conclusion and Future Work
InFL-UX, demonstrated the feasibility of an interactive, browser-based FL system that integrates state-of-the-art technologies from the UI to the server-side aggregation process. This approach allows embedding FL capabilities into IML systems, particularly in privacy-sensitive domains. Despite its limitations, the application effectively utilises ONNX Runtime and modern browser features (WebAssembly and WebGPU) to support client-side FL training, enabling collaborative training across devices while ensuring compatibility and encouraging user adoption.
Future work should involve user studies to assess real-world adoption, performance, and scalability and extend the application to additional computer vision tasks, such as image segmentation and object detection. Moreover, incorporating explainable AI techniques (e.g. class activation maps (Selvaraju et al., 2017), and concept-based modelling (Alam et al., 2025)) could improve transparency and trust. Finally, enabling administrators to deploy new training tasks and allowing clients to participate in multiple tasks with distinct datasets would significantly enhance the system’s utility.
@@ Acknowledgements.
This work was funded, in part, by the German Federal Ministry of Education and Research (BMBF) under grant number 01IW23002 (No-IDLE) and grant number 01IW24006 (NoIDLEChatGPT), by the Lower Saxony Ministry of Science and Culture (MWK) in the zukunft.niedersachsen program, and by the Endowed Chair of Applied AI at the University of Oldenburg.References
@@
- Alam et al. (2025) Hasan Md Tusfiqur Alam, Devansh Srivastav, Md Abdul Kadir, and Daniel Sonntag. 2025. Towards Interpretable Radiology Report Generation via Concept Bottlenecks Using a Multi-agentic RAG. In Advances in Information Retrieval - 47th European Conference on Information Retrieval, ECIR 2025, Lucca, Italy, April 6-10, 2025, Proceedings, Part III (Lecture Notes in Computer Science, Vol. 15574), Claudia Hauff, Craig Macdonald, Dietmar Jannach, Gabriella Kazai, Franco Maria Nardini, Fabio Pinelli, Fabrizio Silvestri, and Nicola Tonellotto (Eds.). Springer, 201–209. https://doi.org/10.1007/978-3-031-88714-7_18
@@
- Tusfiqur et al. (2022) Hasan Md Tusfiqur, Duy MH Nguyen, Mai TN Truong, Triet A Nguyen, Binh T Nguyen, Michael Barz, Hans-Juergen Profitlich, Ngoc TT Than, Ngan Le, Pengtao Xie, et al. 2022. DRG-Net: interactive joint learning of multi-lesion segmentation and classification for diabetic retinopathy grading. arXiv preprint arXiv:2212.14615 (2022).